Find the integral field of the vector field given by X(p)=(p,2(y-b),-2(x-a) through the point (a+r,b) and f(x,y)=(x-a)^2+(y-b)^2 of a circle C=f^-1(r^2)
1
Expert's answer
2020-10-08T16:49:20-0400

1. We point out that the vector field can be represented as "\nabla f=(\frac{\partial f}{\partial p},\frac{\partial f}{\partial x}, \frac{\partial f}{\partial y})" . Indeed, if we integrate the first component with respect to p, we receive: "f=\frac{1}{2}p^2+\alpha(x,y)" . It implies "\alpha_x=2(y-b)", "\alpha_y=-2(x-a)" . It can be easily checked that the latter system has no common solution. In particular, it follows from the fact that "\alpha_{xy}=2" from the first equation and "\alpha_{yx}=-2" from the second one.

Threfore, we correct the statement: "X(p)=(-2(x(p)-a),2(y(p)-b))," where "x(p)=p" , "y(p)=p" . Now we solve the equation: "-2(p-a)=a+r" . The latter implies: "p=\frac{a-r}{2}" . The second equation then yields: "2(\frac{a-r}2-b)=b" . The latter yields "a-r=3b" . Now we will find an integral "\int_{x_0}^{x_1} ds" , where s is given by the curve "X(p)" and "x_0=X(0),x_1=X(\frac{a-r}{2})" . "\int_{x_0}^{x_1}ds=\int_{0}^{\frac{a-r}{2}}|X'(p)|dp=\int_{0}^{\frac{a-r}{2}}\sqrt{4(1-a)^2+4(1-b)^2}dp=\sqrt{4(1-a)^2+4(1-b)^2}\frac{a-r}{2}"

2.As it is stated, the aim is to find "f(C)," where C is a circle with a radius r and center (a,b). We receive an obvious answer: "f(C)=r^2"